This is a pre-algebra question :)
If a person walks 2.4 miles in 1.5 hours, how far can he walk in 3 hours. I appreciate that a tutor help me but i still dont get it. PLz help me. May u go thruough it step by step :).

Found 2 solutions by rfer, solver91311:
Answer by rfer(16322)

2.4/1.5=x/3
1.5x=3*2.4
x=7.2/1.5
x=4.8

Answer by solver91311(24713)



If you can walk a certain distance is a certain amount of time, how far can you walk in twice that amount of time? For example, if I can walk 6 feet in one second, I should be able to walk 12 feet in two seconds, right? So, if 3 hours is two times 1.5 hours, and I can go 2.4 miles in 1.5 hours, how far can I go in two times that long?

John

My calculator said it, I believe it, that settles it